OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (0, 5) and (–2, 11)? please i need this question answered :(

OpenStudy (anonymous):

Find Slope between the points (5-11)/(0 + 2) -6 / 2 =-3

OpenStudy (anonymous):

y – 5 = 3(x + 2) y – 11 = –3(x – 2) y – 5 = –3(x + 2) y – 11 = –3(x + 2) possible answers ^^^

OpenStudy (anonymous):

y- y1 = m (x-x1) y-11=-3(x+2)

OpenStudy (anonymous):

thankssss sooo muchh